OpenStudy (anonymous):

Can someone please help will medal and fan! A. Describe the difference between 0 slope and undefined slope and give an example of each. B. Describe the different uses for the slope intercept form, the point slope form and the standard form and give examples of each. C. Describe how to determine if two lines are parallel, perpendicular or neither and give examples of each.

jimthompson5910 (jim_thompson5910):

if a line has a slope of 0, then the line is horizontal 

jimthompson5910 (jim_thompson5910):

if the slope is undefined, then the line is completely vertical 

OpenStudy (anonymous):

and undefined slopes don't exist right?

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you help me with b and c please

OpenStudy (anonymous):

@jim_thompson5910 r u there?

jimthompson5910 (jim_thompson5910):

sorry the website is being laggy on my end, one moment

OpenStudy (anonymous):

Okay its ok take your time :)

jimthompson5910 (jim_thompson5910):

how far did you get with part b) ?

OpenStudy (anonymous):

oh i don't really understand these three questions thats why i came here for help :)

jimthompson5910 (jim_thompson5910):

do you have in your notes `slope intercept form` ?

OpenStudy (anonymous):

y = mx + b

jimthompson5910 (jim_thompson5910):

what does m and b represent?

OpenStudy (anonymous):

m is the slope and b is the y intercept

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

so if you know the slope and y intercept, you can form the equation or you can go in reverse if you know the equation, you can find the slope and y intercept

OpenStudy (anonymous):

ya

jimthompson5910 (jim_thompson5910):

what is point slope form?

OpenStudy (anonymous):

y - k = m(x - h)

OpenStudy (anonymous):

m is the slope and h,k is a point on the line

jimthompson5910 (jim_thompson5910):

or you can write it as \[\Large y - y_1 = m(x - x_1)\] but you pretty much have it

OpenStudy (anonymous):

oh ok

jimthompson5910 (jim_thompson5910):

your way is probably better/easier to write on a keyboard

OpenStudy (anonymous):

ya

jimthompson5910 (jim_thompson5910):

what is one advantage of point slope form?

OpenStudy (anonymous):

its an easy way of finding a line equation with two points

jimthompson5910 (jim_thompson5910):

yes, you can find the slope of the line through those two points then you can use point slope form to plug in the slope m and one of the points (h,k)

OpenStudy (anonymous):

so thats my answer for b right?

jimthompson5910 (jim_thompson5910):

you just need to go over standard form and then you're done with part b

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

um i think standard form is like 10 to the power of 3 = 1000 right?

jimthompson5910 (jim_thompson5910):

standard form, in terms of a linear equation is this Ax + By = C hopefully that looks familiar?

OpenStudy (anonymous):

oh yea x,y is the point and A B C are the integers

jimthompson5910 (jim_thompson5910):

yes (x,y) is on the line

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

and how do we use standard form?

jimthompson5910 (jim_thompson5910):

standard form is useful for helping to find x and y intercepts you just plug in x = 0 to find the y intercept or plug in y = 0 to find the x intercept

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

and C i'm not sure about because like are they asking about normal lines or...

jimthompson5910 (jim_thompson5910):

if two lines are parallel, what must be true of their slopes?

OpenStudy (anonymous):

its not a function......

OpenStudy (anonymous):

i don't know

jimthompson5910 (jim_thompson5910):

the slopes would be equal

jimthompson5910 (jim_thompson5910):

example y = 2x+7 y = 2x+9 both slopes are 2. If you graphed the two equations, you'll find that they never cross

OpenStudy (anonymous):

ohhh ok

jimthompson5910 (jim_thompson5910):

if two lines are perpendicular, then the two slopes will multiply to -1 example y = 3x+5 y = -1/3x + 10 3 times -1/3 = -1, so these two lines are perpendicular. They meet at a 90 degree angle

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and i think we're done now right?

jimthompson5910 (jim_thompson5910):

yes I think that covers it all

OpenStudy (anonymous):

oh good thank you soo much :)

jimthompson5910 (jim_thompson5910):

no problem